f(x) = x^2+2x and g(x)= 2-x
Find (f-g)(x) and
(f-g)(9)

Answer by ikleyn(52778)

.
f(x) = x^2+2x and g(x)= 2-x

Find (f-g)(x)             (f-g)(x) = x^2 + (2x - x) + 2 = x^2 + x + 2.     Solved. 

(f-g)(9)                  = 9^2 + 9 + 2.                                   Calculate it on your own.